# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v12

@@ B.1 Tsukishima Survey: The First Social Survey in Japan
According to Kawai (1980), the methodology of social research in Japan was formed in the 1930s. During the 1910s and 1920s, “numerous social surveys were conducted by various survey entities, including government agencies, research groups, local administrative bodies, private organizations, survey research institutes, and researchers.” He saw that circa WWI, there were many problems in labor, lifestyle, urban, health and sanitation, public safety, and citizens due to the heavy and chemical industrialization, the monopolization of capitalism, and urban expansion (p.48). The Report on Field Survey in Tsukishima, Kyobashi Ward, Tokyo (called Tsukishima Survey) [^87] Takano himself, who was responsible for overseeing the survey, refers to this field survey as the “Tsukishima Survey” (Department of Health, Ministry of the Interior 1923a, p. 6). Considering this, I use the Tsukishima Survey as the name of this survey throughout this paper. used in this study is the most pioneering survey of urban communities (Sekiya 1970, p. 43).
@@
First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey. The person responsible for the Tsukishima Survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University. Takano became a professor of economics at Tokyo Imperial University in 1919, and became the director of the OISR in 1920. Takano placed great importance on gathering objective statistics about workers through social surveys (Sekiya 1970, p. 33). He had the experience of conducting the first household survey in Japan in 1916888888This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p. 33). The original documents of this survey, preserved at Tokyo Imperial University, were destroyed in the Great Kantō Earthquake (Takano 1933, pp. 727–728). Before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University. The findings were reported in April 1894 as an exercise entitled “East London in Tokyo” (Kawai 1981b, p. 3). He was a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.898989The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No. 172. Initially, 34 members were appointed, and the first meeting was held on July 8. The reason for the establishment of this committee was the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period. Another reason was the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp. 10–13). In 1917, the seventh subcommittee evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp. 12–13). On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions. The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p. 6). Initially, the survey area was planned to be Yanagibashi Yokogawa-chō in Honjo Ward. However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled worker households than in Yokogawa-chō (Kawai 1980, pp. 55–59). [^90] Takeda (2015, pp. 75–76) provides more details on how the survey site was changed. Takano was looking for a commissioned researcher to conduct the survey and asked Sakuzō Yoshino, a colleague at Tokyo Imperial University, to recommend a suitable person. Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position. The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda (2015, pp. 76–77) used the diary to analyze the circumstances surrounding the change of survey location. Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network. On October 26, Tanahashi and Yamana—together with the Kyobashi Branch of the fraternity club—conducted an on-site inspection of Tsukishima. Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima. Finally, Takano and Yamana conducted an on-site inspection there on October 30 and decided on Tsukishima as the survey site.
@@
Although the statistical sampling method was not yet common at that time, Takano selected the survey sites based on his visionary thinking. He stated that “there can be no dispute that we have chosen Tokyo as the target of our metropolitan survey. However, it is impossible to survey the entire Tokyo area. We have no choice but to conduct a partial survey. For this purpose, it would be the best idea to select a representative area in Tokyo and survey that area narrowly but deeply.” (Department of Health, Ministry of the Interior 1923a, pp. 2–3). Takano further explained the need to focus on the working-class, which makes up most of the citizens, rather than to see the political and commercial center: “Most of Tokyo citizens are wage earners who subsist on labor. As is usual in large cities, a large number of people are engaged in manual labor in Tokyo. However, the most genuine type of worker is the skilled worker. Therefore, the area where many families of skilled workers congregate can be regarded as a representative area of Tokyo. Further, it seemed to me that a study of the social and sanitary conditions in that area could explain a large part of the social and sanitary conditions in Tokyo.” (Department of Health, Ministry of the Interior 1923a, pp. 3).
@@
It is necessary to investigate the point where Takano states that “the most genuine type of worker is the skilled workers” and the fact that the main target of the Tsukishima Survey was skilled workers. [^91] In other words, the Tsukishima Survey was not designed for poor households. Therefore, unlike the surveys conducted in the UK by Booth and Rowntree, investigating the standard of living was not the main objective of the survey (Sekiya 1970, p. 42). The description by Kawai (1981a, pp. 13–18) is helpful in this regard. As summarized in Section 2, Japan’s mechanical and chemical industrialization throughout WWI led to a large concentration of city workers. As a result, the ratio of small-scale conventional factories based on “master employees” (as is often observed in the textile industry) decreased. Conversely, the number of factories with a direct “factory owner-employee” employment relationship increased, particularly in the machinery and equipment, metal, and chemical industries. These workers in the modern industrial sector accumulated skills, formed households, and settled in the cities. It can be said that a new urban worker lifestyle was being formed. In short, although the city had a class of workers engaged in traditional industries, Takano believed that the skilled workers in the modern industrial sector would play a central role in terms of the labor force, labor movement, and urban life.
@@
In addition, Sekiya notes the rise of the labor movement. As explained in Section 2, while wages rose throughout WWI, inflation caused relative wages to decline. In 1916, Japan’s first labor protection legislation, the Factory Law, was enacted. However, the number of working days and hours did not decrease. Consequently, the number of labor disputes steadily increased (Sekiya 1970, pp. 24–26). Given this, Takano believed that workers should organize labor unions to help themselves with labor problems that could not be resolved through labor protection legislation. Sekiya (1970, pp. 26–40) speculates that Takano’s intention in conducting the Tsukishima Survey was to obtain basic data for solving labor problems, specifically to identify the scope of matters to be solved by workers on their own initiative (or covered by labor protection legislation).
@@
Kawai (1981b, pp. 15; 17) assessed the $40$ households covered by the household survey included in the Tsukishima Survey as “the representativeness of the sample has remained ambiguous.” There are no systematic descriptions of how the sample households were characterized among the worker households in Tsukishima. The official report stated that “to find suitable persons to fill in the form, we must ask for the assistance of school principals, police officers, physicians, and laborers, and we have requested a meeting of these persons at the survey office to explain the purpose of the survey, and to solicit applicants for the distribution of the household account book” (Department of Health, Ministry of the Interior 1923a, pp. 8–9). As explained, the target households were not randomly selected because the statistical sampling method was not prevalent. However, as described in Section 3, the THBS households can be regarded as standard skilled factory worker households in Tsukishima. Therefore, it is plausible that Takano designed this household budget survey in the sampling framework to capture representative households in Tsukishima.
@@ 10. Elementary school survey July 1919 2
Notes: The years/months listed in the first column (“Periods of the field studies”) are based on the description in the Department of Health, Ministry of the Interior (1923a, pp. 10–17). Figures listed in the second column (“Chapter in the published report”) are based on the table of contents in the Department of Health, Ministry of the Interior (1923a, pp. 1–12; 1923b, pp.1–2). “A2” indicates the Appendix 2.
Sources: Department of Health, Ministry of the Interior (1923a, 1923b, 1923c).
Next, I overview the survey items and methodology. The investigation was primarily divided into “documentary” and “field” surveys. The former was an analysis of the characteristics of Tsukishima, using published statistics such as the Police Department Statistics Book (Department of Health, Ministry of the Interior 1923a, p. 10). This was a pioneering attempt at the time and influenced subsequent urban studies (Kawai 1980, pp. 61–62). The latter was a field survey including $14$ survey items (Table B.1). It is not possible to describe all the survey items in detail, but following the report, they can be divided into four large categories: 1. survey on workers’ livelihood (chapter 2); 2. survey on sanitary conditions (chapter 3); 3. survey on labor conditions (chapter 4); and 4. social mapping and photography (appendix). The survey was conducted over two years, from November 1918 to November 1920, although the timing varied depending on the survey items (Table B.1).
A unique feature of the survey methodology is the use of young researchers as commissioned researchers. The household budget survey and leisure life survey were conducted by Yasunosuke Gonda, a faculty member of the Doitsugaku kyōkai gakkō (Association for German Sciences), who had researched cinematograph and entertainment. Tetsuo Hoshino, a medical doctor, was responsible for the sanitary survey. The labor survey was managed by Yoshitsuru Yamana, a graduate of Tokyo Imperial University who contributed to the organization of labor unions. In addition, Toyotaro Miyoshi, a sociology student at the Tokyo Imperial University, and others became temporary employees (see footnote 54). At the time, the Health Bureau of the Ministry of the Interior conducted several sanitary surveys of rural areas. Takano may have designed the urban sanitary survey from this background (Kawai 1981a, pp. 8–9). However, based on the idea that “health and sanitation surveys can at least encompass economic and social surveys,” Takano included a wide range of survey items, not limited to sanitation surveys (Department of Health, Ministry of the Interior 1923a, p. 56). Kawai (1980, p. 55) credits the Tsukishima Survey as a noteworthy pioneering concept: “going beyond a mere insurance and sanitation survey, it attempted to accurately depict the state of urban workers’ lives as they became established in the city during modernization, with skilled workers as its core.”
@@
Takano stated, “To be as complete as possible, it is best to continue the survey over as long period as possible and to obtain accurate facts by recording the workers’ household income and expenditures.” For this reason, he believed that “the best method of investigation would be to establish a survey station directly in the survey area and to have a full-time investigator reside there as much as possible, familiarize himself with the area and the people, and conduct the field survey” (Department of Health, Ministry of the Interior 1923a, p. 4). In fact, through the mediation of the Tsukishima police station, a survey office was rented at 9-3 Higashinakadōri, Tsukishima, where Yoshitsuru Yamana moved in from November 1918 and contacted worker households (Department of Health, Ministry of the Interior 1923a, p. 7; Takeda 2015, p. 77). The survey office was open from November 1918 to December 1920, including the entire survey period. Figure B.1 shows the photographs of the survey office.
It is also important to note that Takano and his team held weekly meetings at the office during the survey period. Toyotaro Miyoshi, who was an assistant in the survey, recalls the situation at that time as follows:
@@
He also states the following about Takano:
@@
Kawai (1981b, p. 28) assessed that there were no interconnections among the survey items mentioned above and that the survey “ended up being several fragmentary surveys rather than a comprehensive and systematic survey of workers’ lives.” This difficulty may explain the reason why previous academic studies have rarely used the Tsukishima Survey. However, as already explained, Takano did not intend to conduct a large-scale social survey in Tsukishima in the first place. Instead, as Takano noted, the best method is to investigate a representative area “narrowly but deeply.” The goal was to improve the quality of the investigation rather than the scale of it. In this light, the fact that the researchers lived in the survey area and were constantly organizing and revising the survey results should be appreciated. Although the Tsukishima Survey is an early social survey in Japan, it is undoubtedly of high quality.
@@
Notes: This photograph shows an example of the official report of the Tsukishima Survey published by the Department of Health, Ministry of the Interior, in 1923.
Source: Department of Health, Ministry of the Interior 1923b, title page. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
Finally, I summarize the format of the published report and the unpublished materials. The results of the Tsukishima Survey were officially summarized in one report published in December 1921 by the Sanitary Bureau of the Ministry of the Interior and in two appendices (Figure B.2 is an example of the front cover). The report was published under the name “Report on Field Survey in Tsukishima, Kyobashi, Tokyo,” and consists of four sections. The first section is a comprehensive review by Takano and outlines the survey. The second section summarizes the living conditions of the workers, and was written by Gonda (partly by Yamana). Hoshino wrote the third section. It summarizes the sanitary conditions of workers. In the fourth section, Yamana summarizes the labor situation in Tsukishima. There are two appendices. One contains statistical tables related to the above report, and the other organizes social maps and photographs. The correspondence between the survey items and the report is summarized in Table B.1. The Ministry of Home Affairs has already published these materials, which are available at the Gonda Yasunosuke Library of the OISR and the University of Tokyo Library. [^92] In the OISR, these are included in the Archives of Gonda Yasunosuke (7-2; 7-3; 7-4). In addition, there is a second manuscript of the Tsukishima Survey report. This is the manuscript of the second series, which was preserved in the home of Gonda Hayao, the second son of Gonda Yasunosuke, and was discovered by Terade Koji. The second series includes a summary of the results of a household survey of elementary school teachers’ households in Tokyo, conducted as a supplementary survey to the Tsukishima Survey (see Terade (1982) for details). The manuscript of this second series was later donated to the OISR and is now available as part of the Gonda Yasunosuke Collection at the Institute (The OISR, Archives of Gonda Yasunosuke (8-12; 8-13; 8-14; 8-15)).

### v14

@@ B.1 Tsukishima Survey: The First Social Survey in Japan
According to Kawai (1980), the methodology of social research in Japan was established in the 1930s. During the 1910s and 1920s, “numerous social surveys were conducted by various entities, including government agencies, research groups, local administrative bodies, private organizations, survey research institutes, and individual researchers.” The Report on Field Survey in Tsukishima, Kyobashi Ward, Tokyo (called the Tsukishima Survey) represents a pioneering survey of urban communities (Sekiya 1970, p. 43).
@@
First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey. [^71] Takano himself refers to this field survey as the “Tsukishima Survey” (Department of Health, Ministry of the Interior 1923a, p. 6). Considering this, I use the Tsukishima Survey as the name of this survey throughout this paper. The person responsible for the survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University. Circa WWI, Takano observed widespread problems in labor, lifestyle, urban infrastructure, health and sanitation, public safety, and citizen welfare, driven by heavy and chemical industrialization, capitalist monopolization, and urban expansion. In this context, Takano emphasized the importance of gathering objective statistics about workers through social surveys (Sekiya 1970, p. 33). He conducted Japan’s first household survey [^72] This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p. 33). The original documents of this survey, preserved at Tokyo Imperial University, were destroyed in the Great Kantō Earthquake (Takano 1933, pp. 727–728). Before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University. The findings were reported in April 1894 as an exercise entitled “East London in Tokyo” (Kawai 1981b, p. 3). and had served on the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.737373The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No. 172. Initially, 34 members were appointed, and the first meeting was held on July 8. The reason for the establishment of this committee was the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period. Another reason was the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp. 10–13). In 1917, the seventh subcommittee evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp. 12–13). On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions. The survey aimed to determine the housing conditions, household budgets, and health status (including child health, live births, stillbirths, deaths, and diseases) in areas where skilled-worker families resided (Sekiya 1970, p. 6). Initially, the survey area was planned as Yanagibashi Yokogawa-chō in Honjo Ward. However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled-worker households than Yokogawa-chō (Kawai 1980, pp. 55–59). [^74] Takeda (2015, pp. 75–76) provides more details on how the survey site was changed. Takano was looking for a commissioned researcher to conduct the survey and asked Sakuzō Yoshino, a colleague at Tokyo Imperial University, to recommend a suitable person. Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position. The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda (2015, pp. 76–77) used the diary to analyze the circumstances surrounding the change of survey location. Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network. On October 26, Tanahashi and Yamana—together with the Kyobashi Branch of the fraternity club—conducted an on-site inspection of Tsukishima. Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima. Finally, Takano and Yamana conducted an on-site inspection there on October 30 and decided on Tsukishima as the survey site.
@@
Although statistical sampling was not yet common at the time, Takano selected the survey site based on his judgment. He stated: “There can be no dispute that we have chosen Tokyo as the target of our metropolitan survey. However, it is impossible to survey the entire Tokyo area. We have no choice but to conduct a partial survey. For this purpose, it would be the best idea to select a representative area in Tokyo and survey that area narrowly but deeply” (Department of Health, Ministry of the Interior 1923a, pp. 2–3). Takano further emphasized the importance of focusing on the working class, which constitutes the majority of citizens, rather than the political and commercial center: “Most of Tokyo citizens are wage earners who subsist on labor. As is usual in large cities, a large number of people are engaged in manual labor in Tokyo. However, the most genuine type of worker is the skilled worker. Therefore, the area where many families of skilled workers congregate can be regarded as a representative area of Tokyo. Further, it seemed to me that a study of the social and sanitary conditions in that area could explain a large part of the social and sanitary conditions in Tokyo” (Department of Health, Ministry of the Interior 1923a, pp. 3).
@@
Examining Takano’s statement that “the most genuine type of worker is the skilled worker” and the fact that the main target of the Tsukishima Survey was skilled workers is important. [^75] In other words, the Tsukishima Survey was not designed for poor households. Therefore, unlike the surveys conducted in the UK by Booth and Rowntree, investigating the standard of living was not the main objective of the survey (Sekiya 1970, p. 42). Kawai’s description (1981a, pp. 13–18) is helpful in this regard. As summarized in Section 2, Japan’s mechanical and chemical industrialization during WWI led to a large concentration of city workers. Consequently, the ratio of small-scale conventional factories based on “master employees” (as commonly seen in the textile industry) declined. By contrast, the number of factories with a direct “factory owner–employee” relationship increased, particularly in the machinery and equipment, metal, and chemical industries. Workers in the modern industrial sector acquired skills, formed households, and settled in cities, giving rise to a new urban worker lifestyle. Although the city still had a class of workers engaged in traditional industries, Takano considered skilled workers in the modern industrial sector to be central in terms of labor force, labor movements, and urban life.
@@
Sekiya also notes the rise of the labor movement. As explained in Section 2, although wages rose during WWI, inflation caused relative wages to decline. In 1916, Japan enacted its first labor protection legislation, the Factory Law. However, the number of working days and hours remained unchanged, leading to a steady increase in labor disputes (Sekiya 1970, pp. 24–26). Takano therefore believed that workers should organize labor unions to address problems not resolved through labor protection legislation. Sekiya (1970, pp. 26–40) speculates that Takano’s purpose in conducting the Tsukishima Survey was to obtain basic data for solving labor issues, specifically to identify problems that workers could address independently or through existing labor legislation.
@@
Kawai (1981b, pp. 15; 17) assesses the $40$ households covered by the Tsukishima Survey as having “ambiguous representativeness.” No systematic description exists of how sample households were characterized among worker households in Tsukishima. The official report noted: “To find suitable persons to fill in the form, we must ask for the assistance of school principals, police officers, physicians, and laborers, and we have requested a meeting of these persons at the survey office to explain the purpose of the survey, and to solicit applicants for the distribution of the household account book” (Department of Health, Ministry of the Interior 1923a, pp. 8–9). As noted, the target households were not randomly selected because statistical sampling was not yet prevalent. However, as described in Section 3, the THBS households can be regarded as standard skilled factory-worker households in Tsukishima. Therefore, Takano may have designed this household budget survey to capture representative households within the sampling framework of Tsukishima.
@@ 10. Elementary school survey July 1919 2
Notes: The years/months listed in the first column (“Periods of the field studies”) are based on the description in the Department of Health, Ministry of the Interior (1923a, pp. 10–17). Figures listed in the second column (“Chapter in the published report”) are based on the table of contents in the Department of Health, Ministry of the Interior (1923a, pp. 1–12; 1923b, pp.1–2). “A2” indicates the Appendix 2. Sources: Department of Health, Ministry of the Interior (1923a, 1923b, 1923c).
Next, I provide an overview of the survey items and methodology. The investigation was primarily divided into “documentary” and “field” surveys. The former analyzed the characteristics of Tsukishima using published statistics, such as the Police Department Statistics Book (Department of Health, Ministry of the Interior 1923a, p. 10). This was a pioneering approach that influenced subsequent urban studies (Kawai 1980, pp. 61–62). The field survey included $14$ items (Table B.1). According to the report, these can be grouped into four main categories: 1. workers’ livelihood (chapter 2); 2. sanitary conditions (chapter 3); 3. labor conditions (chapter 4); and 4. social mapping and photography (appendix). The survey was conducted over two years, from November 1918 to November 1920, with timing varying depending on the survey items (Table B.1).
A distinctive feature of the methodology was the use of young commissioned researchers. The household budget and leisure life surveys were conducted by Yasunosuke Gonda, a faculty member of the Doitsugaku Kyōkai Gakkō (Association for German Sciences) who had conducted research on cinematography and entertainment. Tetsuo Hoshino, a medical doctor, oversaw the sanitary survey. Yoshitsuru Yamana, a Tokyo Imperial University graduate who contributed to labor union organization, managed the labor survey. Additionally, Toyotaro Miyoshi, a sociology student at Tokyo Imperial University, and others served as temporary employees (footnote 40). At the time, the Health Bureau of the Ministry of the Interior conducted several rural sanitary surveys, and Takano may have drawn on this experience in designing the urban sanitary survey (Kawai 1981a, pp. 8–9). However, operating on the principle that “health and sanitation surveys can at least encompass economic and social surveys,” Takano included a broad range of survey items beyond sanitation (Department of Health, Ministry of the Interior 1923a, p. 56). Kawai (1980, p. 55) credits the Tsukishima Survey as a notable pioneering effort: “Going beyond a mere insurance and sanitation survey, it attempted to accurately depict the state of urban workers’ lives as they became established in the city during modernization, with skilled workers as its core.”
@@
Takano stated, “To be as complete as possible, it is best to continue the survey over as long a period as possible and to obtain accurate facts by recording the workers’ household income and expenditures.” For this reason, he believed that “the best method of investigation would be to establish a survey station directly in the survey area and to have a full-time investigator reside there as much as possible, familiarize himself with the area and the people, and conduct the field survey” (Department of Health, Ministry of the Interior 1923a, p. 4). With the mediation of the Tsukishima police station, a survey office was rented at 9-3 Higashinakadōri, Tsukishima. Yoshitsuru Yamana moved in from November 1918 and began contacting worker households (Department of Health, Ministry of the Interior 1923a, p. 7; Takeda 2015, p. 77). The office remained open for the entire survey period, from November 1918 to December 1920. Figure B.1 shows photographs of the survey office.
Importantly, Takano and his team held weekly meetings at the office throughout the survey period. Toyotaro Miyoshi, who worked as an assistant on the survey, recalls:
@@
He also notes:
@@
Kawai (1981b, p. 28) observes no clear interconnections among the survey items, and that the survey “ended up being several fragmentary surveys rather than a comprehensive and systematic survey of workers’ lives.” This may explain why previous academic studies have rarely used the Tsukishima Survey. However, as explained earlier, Takano did not aim to conduct a large-scale social survey in Tsukishima. Instead, he argued that the best method is to investigate a representative area “narrowly but deeply,” prioritizing the quality of the investigation over its scale. In this light, the fact that researchers lived in the survey area and continuously organized and revised the results is significant. Although the Tsukishima Survey is an early social survey in Japan, it is undoubtedly of high quality.
@@
Notes: This photograph shows an example of the official report of the Tsukishima Survey published by the Department of Health, Ministry of the Interior, in 1923. Source: Department of Health, Ministry of the Interior 1923b, title page. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
Finally, I summarize the format of the published report and the unpublished materials. The results of the Tsukishima Survey are officially documented in a report published in December 1921 by the Sanitary Bureau of the Ministry of the Interior, along with two appendices (Figure B.2 shows an example of the front cover). The report, published under the title “Report on Field Survey in Tsukishima, Kyobashi, Tokyo,” consists of four sections. The first section, a comprehensive review by Takano, outlines the survey. The second section, written by Gonda (partly by Yamana), summarizes the living conditions of workers. Hoshino authored the third section, which addresses workers’ sanitary conditions. In the fourth section, Yamana summarizes the labor situation in Tsukishima. The two appendices complement the main report. One contains statistical tables corresponding to the report, while the other organizes social maps and photographs. The correspondence between survey items and the report is summarized in Table B.1. These materials have been published by the Ministry of Home Affairs and are available at the Gonda Yasunosuke Library of the OISR and the University of Tokyo Library. At the OISR, they are included in the Archives of Gonda Yasunosuke (7-2; 7-3; 7-4).